OpenStudy (anonymous):

Help please!? this one is so confusing! Simplify (x + 2)(x2– 4x – 9) x3– 2x2– 17x – 18 x3– 6x2– x – 18 x3– 4x2– 3x – 18 x3– 10x2– 9x – 18 THANK YOU! ::)

OpenStudy (anonymous):

also this one. What does it mean factor completely? but I'm guessing its D? Factor completely 3x3– 21x2– 27x 3x(x2– 7x – 9) 3x(x + 1)(x – 9) 3(x2– 7x – 9) 3(x3– 21x2– 27x)

OpenStudy (anonymous):

answer is x3– 2x2– 17x – 18

OpenStudy (anonymous):

wait no...C for the seccond Q

OpenStudy (anonymous):

(x + 2)(x2– 4x – 9) = x^3 -4x^2 - 9x+2x^2 - 8x -18= x^3 -2x^2 -17x -18...so option A

OpenStudy (anonymous):

For secnd Q it is option A...pretty dure

OpenStudy (anonymous):

3x3– 21x2– 27x = 3x( x^2 -7x - 9)...taking 3x common...